This came from Factoring polynomials Section
Solve..
36=1/2x+(x+4)

Answer by stanbon(75887)

Solve..
36=1/2x+(x+4)
---------------------
multiply thru by 2x to get:
72x = 1 + 2x(x+4)
72x = 1 + 2x^2 + 8x
2x^2 - 64x +1 = 0
x = [64 +- sqrt(64^2-4*2*1)]/4
x = [64 +- sqrt(4089)]/4
x = [64 +- 63.945]/4
x = 31.9863 = x = 0.01368...
--------------------------------------------
Cheers,
Stan H.